\begin{thebibliography}{99}
\bibitem{Gouy1} CR Gouy. Sur une proprieté nouvelle des ondes lumineuses. C R Acad Sci Paris 110, 1251 (1890).
\bibitem{Gouy2} CR Gouy. Sur la propagation anomale des ondes. Ann Chim Phys Ser 6, 24, 145 (1891).
\bibitem{Michigan} S Feng and HG Winful. Physical origin of the Gouy phase shift. Opt Lett 26, 485 (2001).
\bibitem{Michi2} S Feng, HG Winful, and RW Hellwarth. Gouy shift and temporal reshaping of focused single-cycle electromagnetic pulses. Opt Lett 23, 385 (1998).
\bibitem{Siegman} AE Siegman. Lasers. University Science, Mill Valley, Calif. 1986.
\bibitem{Carmelo} RY Zhong, ZH Zhu, HJ Wu, C Rosales-Gusmán, SW Song and BS Shi. Gouy-phase-mediated propagation variations and revivals of transverse structure in vectorially structured light. Phys Rev A 103, 053520 (2021).
\bibitem{Photonics} JL Jia, KP Zhang, GWHu, MP Hu, T Tong, QQ Mu, H Gao, FL Li, CW Qiu and P Zhang. Arbitrary cylindrical vector beam generation enabled by polarization-selective Gouy phase shifter. Photonic Res 9, 1048-1054 (2021).
\bibitem{Ducharme} RJ Ducharme, IG da Paz, and AG Hayrapetyan. Fractional Angular Momenta, Gouy and Berry Phases in Relativistic Bateman-Hillion-Gaussian Beams of Electrons. Phys Rev Lett 126, 134803 (2021).
\bibitem{JOSAA} S Pelchat-Voyer and M Piche. Clarifications on the Gouy phase of radially polarized laser beams. J Opt Soc Am A 37, 1496-1504 (2020).
\bibitem{ACS}  T Lee, Y Cheong, HW Baac and LJ Guo, Origin of Gouy phase shift identified by laser-generated focused ultrasound. ACS Photonics 7, 3236-3245 (2020).
\bibitem{Sivil} GA Siviloglou, J Broky, A Dogariu, DN Christodoulides. Observation of accelerating airy beams. Physical Review Letters 99 (21) (nov 2007).
\bibitem{Vo} S Vo, K Fuerschbach, KP Thompson, MA Alonso, JP Rolland. Airy beams: a geometric optics perspective. Journal of the Optical Society of America A 27 (12) (2010) 2574.
\bibitem{Efrem} NK Efremidis, Z Chen, M Segev, DN Christodoulides. Airy beams and accelerating waves: an overview of recent advances. Optica 6 (5) (2019) 686.
\bibitem{Cerda} S Chávez-Cerda, U Ruiz, V Arrizón, HM Moya-Cessa. Generation of airy solitary-like wave beams by acceleration control in inhomogeneous media. Optics Express 19 (17) (2011) 16448.
\bibitem{Jaure} R Jáuregui, PA Quinto-Su. On the general properties of symmetric incomplete airy beams. Journal of the Optical Society of America A 31 (11) (2014) 2484.
\bibitem{Torre} A Torre. Propagating airy wavelet-related patterns. Journal of Optics 17 (7) (2015) 075604.
\bibitem{ChileI} FA Asenjo, SA Hojman, HM Moya-Cessa and F Soto-Eguibar.  Propagation of light in linear and quadratic GRIN media: The Bohm potential.Optics Communications 490, 126947 (2021).
\bibitem{ChileII} SA Hojman, FA Asenjo, HM Moya-Cessa and F Soto-Eguibar. Bohm potential is real and its effects are measurable.
Optik 232, 166341 (2021).
\bibitem{Hojman} SA Hojman and FA Asenjo. A new approach to solve the one-dimensional Schrödinger equation using a wavefunction potential. Phys Lett A 384 (2020) 126913.
\bibitem{Hojman2} SA Hojman and FA Asenjo. Quantum particles that behave as free classical particles. Phys Rev A 102 (2020) 052211.
\bibitem{Hojman3} SA Hojman and FA Asenjo. Dual wavefunctions in two-dimensional quantum mechanics. Phys Lett A 384 (2020) 126263.
\bibitem{makowski} AJ Makowski and S Konkel. Identical motion in classical and quantum mechanics. Phys Rev A 58, (1998) 4975-4977.
\bibitem{Madelung} E Madelung. Quantentheorie in hydrodynamischer form. Z Phys 40 (1927) 322-326.
\bibitem{Bohm} D Bohm. A Suggested Interpretation of the Quantum Theory in Terms of "Hidden" Variables. I. Phys Rev 85 (1952) 166.
\bibitem{Lewis0}  HR Lewis. Classical and Quantum Systems with Time-Dependent Harmonic-Oscillator-Type Hamiltonians. Phys Rev Lett (1967) 18, 510-513.
\bibitem{Lewis} HR Lewis and  PGL Leach. J Math Phys (1982) 23, 165-175.
\bibitem{Ray} JR Ray. Exact solutions to the time-dependent Schrödinger equation Phys Rev A (1982), 26, 729-733.
\bibitem{Thylwe} KE Thylwe and  HJ Korsch. The 'Ermakov-Lewis' invariants for coupled linear oscillators J Phys (1998), 31,  L279-L285.
\bibitem{JPA} M Fernández-Guasti and HM Moya-Cessa. Amplitude and phase representation of quantum invariants for the time dependent harmonic oscillator.  Physical Review A 67 (6), 063803 (2003).
\bibitem{Moya2007} H Moya-Cessa and  M Fernández-Guasti. Time dependent harmonic oscillator subject to a sudden change of mass: continuous solution. Rev Mex  Fis (2007) 53 42-6.
\bibitem{Pedrosa} IA Pedrosa. Exact wave functions of a harmonic oscillator with time-dependent mass and frequency. Phys Rev A 1997 55 3219.
\bibitem{Riesen} HR Lewis and WB Riesenfeld. An Exact Quantum Theory of the Time-Dependent Harmonic Oscillator and of a Charged Particle in a Time-Dependent Electromagnetic Field. J Math Phys 1969 101458.
\bibitem{ZelayaI} K Zelaya and O Rosas-Ortiz. Quantum nonstationary oscillators: Invariants, dynamical algebras and coherent states via point transformations. Phys. Scr. 95, 064004 (2020).
\bibitem{ZelayaII} S Cruz, R Razo, O Rosas-Ortiz and K Zelaya. Coherent states for exactly solvable time-dependent oscillators generated by Darboux transformations. Phys. Scr. 95 044009 (2020).
\bibitem{ZelayaIII} K Zelaya and V Hussin. Time-dependent rational extensions of the parametric oscillator: quantum invariants and the factorization method. J Phys A: Math Theor 53 165301 (2020).
\bibitem{Marte} MAM Marte and S Stenholm. Paraxial light and atom optics: The optical Schrödinger equation and beyond. Phys Rev A 56, 2940-2953 (1997).
\bibitem{Feynman} RP Feynman. Space-Time Approach to Non-Relativistic Quantum Mechanics. Rev Mod Phys 20, 367 (1948).
\bibitem{PRL} R Keil, A Perez-Leija, F Dreisow, M Heinrich, H Moya-Cessa, S Nolte, DN Christodoulides and A Szameit. Classical analogue of displaced Fock states and quantum correlations in Glauber-Fock photonic lattices. Physical Review Letters 107, 103601 (2011).
\bibitem{Optica} T Eichelkraut, C Vetter, A Perez-Leija, H Moya-Cessa, DN Christodoulides and A Szameit. Coherent random walks in free space. Optica 1, 268-271 (2014).
\bibitem{Kapitza} I Ramos-Prieto, K Uriostegui, J Récamier, F Soto-Eguibar and HM Moya-Cessa. Kapitza-Dirac photonic lattices. Opt Lett 46, 4690-4693 (2021).
\bibitem{Wyatt} RE Wyatt. Quantum Dynamics with Trajectories: Introduction to Quantum Hydrodynamics. Springer, 2005.
\bibitem{Holland}PR Holland. The Quantum Theory of Motion: An Account of the de Broglie-Bohm Causal Interpretation of Quantum Mechanics. Cambridge University Press, 1993.
\bibitem{ahms}  FA Asenjo, SA Hojman, HM Moya-Cessa, and F Soto-Eguibar. Propagation of light in linear and quadratic GRIN media: The Bohm potential. Optics Communications 490, 126947 (2021).
\bibitem{Yuen} HP Yuen. Two-photon coherent states of the radiation field. Phys Rev A  13, 2226-2243 (1976).
\bibitem{Caves}  CM Caves. Quantum-mechanical noise in an interferometer.  Phys Rev D 23, 1693-1708 (1981).
\bibitem{Barnett} SM Barnett, A Beige, A Ekert et al. Journeys from Quantum Optics to Quantum Technology. Progr  Quant Electr 54, 19--45 (2017).
\end{thebibliography}

\documentclass[a4paper,fleqn]{cas-dc}

\usepackage{braket}
\usepackage{hyperref}
\usepackage[numbers]{natbib}
\def\tsc#1{\csdef{#1}{\textsc{\lowercase{#1}}\xspace}}
\tsc{WGM}
\tsc{QE}
\tsc{EP}
\tsc{PMS}
\tsc{BEC}
\tsc{DE}
\begin{document}
\let\WriteBookmarks\relax
\def\floatpagepagefraction{1}
\def\textpagefraction{.001}
\title [mode = title]{Bohm approach to the Gouy phase shift}                      
\author[1]{H\'ector M. Moya-Cessa}
\author[2,3]{Sergio A. Hojman}
\author[3]{Felipe A. Asenjo}
\author[1]{Francisco Soto-Eguibar}
\cormark[1]
\address[1]{Instituto Nacional de Astrofísica Óptica y Electrónica, Calle Luis Enrique Erro No. 1, Santa María Tonantzintla, Pue., 72840, Mexico}
\address[2]{Departamento de Ciencias, Facultad de Artes Liberales, Universidad Adolfo Ibáñez, Santiago 7491169, Chile.\\
Departamento de Física, Facultad de Ciencias, Universidad de Chile, Santiago 7800003, Chile.\\
Centro de Recursos Educativos Avanzados, CREA, Santiago 7500018, Chile.}
\address[3]{Facultad de Ingeniería y Ciencias,
Universidad Adolfo Ibáñez, Santiago 7491169, Chile.}
\cortext[cor1]{feguibar@inaoep.mx}

\begin{abstract}
By adapting the Madelung-Bohm formalism to paraxial wave propagation we show, by using Ermakov-Lewis techniques, that the Gouy phase is related to the form of the phase chosen in order to produce a Gaussian function as a propagated field. For this, we introduce a quantum mechanical invariant, that it is explicitly time dependent despite the fact that the Hamiltonian is itself time-independent. We finally show that the effective Bohm {\it index of refraction} generates a GRIN medium that produces the focusing needed for the Gouy phase.
\end{abstract}

\begin{keywords}
Gouy phase \sep Bohm potential \sep Lewis-Ermakov invariant
\end{keywords}

\maketitle

\section{Introduction}
The Gouy phase \cite{Gouy1,Gouy2,Michigan,Michi2,Siegman} is a phase gradually acquired by a beam around the focal region and it results in an increase in the apparent wavelength near the waist ($z \approx 0$) deriving in the fact that the phase velocity in such region formally exceeds the speed of light. It has been explicitly shown that the Gouy phase shift of any focused beam originates from transverse spatial confinement, that introduces a spread in the transverse momentum and therefore a shift in the expectation value of the axial propagation constant. {\color{black}The simultaneous effect of confinement in space combined with spread in momentum is, of course, reminiscent of Heisenberg's uncertainty principle.} \cite{Michigan}. The Gouy phase has been used to  generate arbitrary cylindrical vector beams \cite{Photonics}, to shape vectorially structured light with custom propagation-evolution properties \cite{Carmelo} and to give a new Bateman-Hillion solution to the Dirac equation for a relativistic Gaussian electron beam \cite{Ducharme}, among other applications. Moreover, its origin has been investigated over the years \cite{Michigan,JOSAA,ACS}.

On the other hand, it is well known that during free propagation light  bends when an initial ($z=0$) Airy field is considered \cite{Sivil,Vo,Efrem,Cerda,Jaure,Torre}, which may be explained by using Madelung-Bohm theory \cite{ChileI,ChileII}. This formalism has been applied to solve  the Schrödinger equation for different systems by taking advantage of their non-vanishing Bohm potentials \cite{Hojman,Hojman2,Hojman3,makowski}. We therefore allow to ask if Madelung-Bohm theory \cite{Madelung,Bohm} could be a way to explain Gouy's phase. We show that indeed it may be explained and to this end we use Ermakov-Lewis techniques in order to use the invariant introduced by Lewis for a time dependent harmonic oscillator \cite{Lewis0,Lewis}, and thus use it for a free particle. 

The manuscript is organized as follows: In the next Section, we show that for a free particle an invariant of the Ermakov-Lewis form may be written which, remarkably, is explicitly time dependent. We then give a solution to the Ermakov equation for this particular case. In Section 3, by taking advantage of the similarity between the Schrödinger equation and the paraxial wave equation, we translate the Madelung-Bohm theory to classical optics. In Section 4, we give an operator solution to the (optical) Bohm equations and propose a Gaussian initial field. By using the Ermakov solution obtained in Section 1, we calculate the Gouy phase. Section 5 is left for conclusions.

\section{Ermakov-Lewis invariant}
In the sixties, Lewis \cite{Lewis0,Lewis} introduced an invariant quantity that has the form (we have set $\hbar=1$ and in this Section dot means derivative with respect to time)
\begin{equation}
 {I}=\frac{1}{2}\left[\frac{ {x}^2}{\rho^2}+\left(\rho  {p}-\dot{\rho} {x}\right)^2\right], \label{1Dinvar}
\end{equation}
with  $\rho$ an auxiliary function that obeys the Ermakov equation \cite{Ray,Thylwe,JPA,Moya2007,Pedrosa}
\begin{eqnarray}\label{Erma}
\ddot{\rho}+\Omega^2(t)\rho=\frac{1}{\rho^3};
\end{eqnarray}
consequently (\ref{1Dinvar}) takes the name Ermakov-Lewis invariant. The operator given in \eqref{1Dinvar} is invariant in the sense that
\begin{eqnarray}
\frac{dI}{dt}=\frac{\partial {I}}{\partial t}+i[  {H},I]=0. 
\end{eqnarray}
It is interesting that even though the time-dependent frequency is zero, i.e. a free particle, the Lewis-Ermakov invariant is maintained for such system. In this case the Ermakov equation reduces to
\begin{eqnarray}\label{ErmaF0}
\ddot{\rho}=\frac{1}{\rho^3},
\end{eqnarray}
and the Hamiltonian is that of the free particle (with mass equal to one), i.e. $H_\text{free}=p^2/2$. Therefore $H_\text{free}$ is not the only invariant for a particle freely evolving but also the explicitly time dependent invariant (\ref{1Dinvar}). From the Lewis-Riesenfeld approach \cite{Riesen}, it follows that the quantum invariant $I$ solves an eigenvalue equation with time-independent eigenvalues and  nonstationary eigenfunctions. Such eigenfunctions are not the solutions of the Schr\"odinger equation, however they may be used to construct such solutions \cite{Riesen}.

For the free particle, the auxiliary Ermakov function takes the form 
\begin{equation}\label{ermasol}
    \rho(t)=(at^2+bt+c)^{\frac{1}{2}},
\end{equation}
that is a solution of equation (\ref{ErmaF0}); \textcolor{black}{with $a,\,b,\,c$ are arbitrary real numbers,} with the constraint $ac-b^2/4=1$. It is worth to mention that for the singular oscillator with constant frequency, there also exists a generalization of the Ermakov invariant which is also explicitly time dependent \cite{ZelayaI,ZelayaII,ZelayaIII}.

\section{Madelung-Bohm approach to paraxial wave propagation}
The optical paraxial propagation equation in one dimension is given by \cite{Marte}
\begin{equation}\label{SE}
i\frac{\partial E(x,z)}{\partial z}= -\frac{1}{2k}\frac{\partial^2 E(x,z)}{\partial x^2}-\frac{1}{2} n^2(x,z)E(x,z),
\end{equation}
where $E(x,z)$ is the propagated field, $k$ is the wave number and $n(x,z)$ is the index of refraction. This equation is formally equivalent to the Schrödinger equation by doing $z\rightarrow t$, $k\rightarrow m$ {\color{black} and $-\frac{1}{2} n^2 \rightarrow V(x,z)$, with $m$ the particle mass, $V(x,z)$ {\it quantum} potential and $n$ the refractive index}. We may paraphrase Feynman \cite{Feynman}, stating that there is always the hope that this point of view will inspire more optical simulations of quantum systems \cite{PRL,Optica,Kapitza}. We may give a solution in terms of a polar decomposition \cite{Hojman,Wyatt,Holland}
\begin{equation}\label{psi}
E(x,z)=A(x,z)e^{i S(x,z)},
\end{equation}
with $A(x,z)$ and $S(x,z)$ real functions that depend on the propagation distance and the position. We may separate the real and imaginary parts that come from the substitution of (\ref{psi}) in (\ref{SE}); the first equation reading \cite{ahms} 
\begin{equation} \label{H-J}
\frac{1}{2k}S'^2-\frac{1}{2}({n_B}^2+n^2)+\dot{S}=0,
\end{equation}
and the second one, the continuity (probability conservation) equation,
\begin{equation}\label{Prob} 
\frac{1}{2k}(2A'S'+AS'')+\dot{A}=0,
\end{equation}
with the Bohm index of refraction defined by \cite{Madelung,Bohm}
\begin{equation}\label{bohmpot}
-\frac{1}{2}{n_B}^2=-\frac{1}{2k}\frac{A''}{A},
\end{equation}
where the dot now represents the  propagation distance derivative and the prime the space derivatives.

\section{Operator  solution of the continuity equation}
It is possible to rewrite Eq.~\eqref{Prob} as a Schrodinger-like equation; for this, we do
\begin{equation}\label{Operator}
\frac{\partial A}{\partial z}=-\frac{1}{2k}\left(2S'\frac{\partial }{\partial x}+S''\right)A,
\end{equation}
that, by using the  operator $\hat{p}=-i\frac{\partial }{\partial x}$,  may be taken to the form
\begin{equation}\label{OperatorI}
\frac{\partial A}{\partial z}=-\frac{1}{2k}\left(i2S'\hat{p}+S''\right)A.
\end{equation}
By choosing 
\begin{equation}\label{ese}
    S(x,z)=Q(x)\dot{\nu}(z)+\mu(z),
\end{equation} 
\textcolor{black}{where $Q, \, \nu$ and $\mu$ are arbitrary well behaved real functions,} such that $S'=Q'\dot{\nu}$, and using the property $[f(x),\hat{p}]=if'(x)$ to rearrange terms, we have
\begin{equation}
2S'\hat{p}=\dot{\nu}(Q'\hat{p}+Q'\hat{p})=\dot{\nu}[Q'\hat{p}+\hat{p}Q'+iQ''];
\end{equation}
so that, we may write Eq. (\ref{OperatorI}) as
\begin{equation}
\frac{\partial A}{\partial z}=-i\frac{\dot{\nu}}{2k}\left(Q'\hat{p}+\hat{p}Q'\right)A.
\end{equation}
The function $\mu(z)$ in equation (\ref{ese}) will define the Gouy phase.
The above equation is readily solvable, with  solution
\begin{equation}
A(x,z)=\exp\left\lbrace -\frac{i}{2k}\int\dot{\nu}(z)dz 
\left[Q'(x)\hat{p}+\hat{p}Q'(x)\right] \right\rbrace A_0(x),
\end{equation}
where $A_0(x)=A(x,t=0)$, the initial condition, is an arbitrary (square integrable) function of position.

\subsection{Propagation of a Gaussian field}
Next, we  assume $Q(x)=x^2/2$ and $k=1$ to find the solution
\begin{equation}\label{0110}
A(x,z)=\exp\left[-i\frac{\nu(z)}{2}\left( x\hat{p}+\hat{p}x\right) \right]A_0(x).
\end{equation}
In the above equation, the operator $\exp\left[ -i\frac{\nu(z)}{2}\left( x\hat{p}+\hat{p}x\right)\right] $
is the so-called squeeze operator \cite{Yuen,Caves,Barnett}. By choosing at $z=0$ the amplitude $A_0(x)=\pi^{-1/4} \exp\left(-x^2/2 \right)$, we obtain after application of the squeeze operator
\begin{equation}
A(x,z)=\frac{1}{\pi^{1/4}}\exp\left[-\frac{x^2}{2}e^{-2\nu(z)}-\frac{\nu(z)}{2}\right],
\end{equation}
where we have used that 
\begin{eqnarray}
    e^{-i\frac{\nu(z)}{2}\left( x\hat{p}+\hat{p}x\right) }xe^{i\frac{\nu(z)}{2}\left( x\hat{p}+\hat{p}x\right) }=x\exp[-\nu(z)],
\end{eqnarray}
that may be easily found from the Hadamard formula that states that for two operators $\hat{A}$ and $\hat{B}$, $e^{r\hat{A}}\hat{B}e^{-r\hat{A}}=\hat{B}+r[\hat{A},\hat{B}]+\frac{r^2}{2!}[\hat{A},[\hat{A},\hat{B}]]+\dots$.\\
Now, we are ready to calculate the Bohm index of refraction, for which we need $A'(x,z)=-xe^{-2\nu(z)}A(x,z)$, so that 
\begin{equation}
A''(x,z)=\left[{x^2e^{-4\nu(z)}}-{e^{-2\nu(z)}}\right]A(x,z),
\end{equation}
to obtain
\begin{equation}\label{0130}
-\frac{1}{2}{n_B}^2(x,z)=-\frac{x^2}{2}\exp\left[-4\nu(z)\right]+\frac{1}{2}\exp\left[-2\nu(z)\right] .
\end{equation}
Equation (\ref{Prob}) implies that $\dot{\mu}=-\exp\left[-2\nu(z)\right] /2$, such that we obtain from  equation (\ref{H-J})
\begin{equation}\label{0140}
-\frac{1}{2}n^2(x,z)+\frac{x^2}{2}\left(\ddot{\nu}+\dot{\nu}^2-e^{-4\nu}\right)=0.
\end{equation}
We change variables as follows: $\rho=\exp\left(\nu\right) $, thus $\dot{\mu}=-1/\left(2\rho^2\right)$, $\dot{\nu}=\dot{\rho}/\rho$ and $\ddot{\nu}=\left( \rho\ddot{\rho}-\dot{\rho}^2\right)/2 $, that produces
\begin{equation}\label{0150}
-\frac{1}{2}n^2(x,z)+\frac{x^2}{2\rho}\left(\ddot{\rho}-\frac{1}{\rho^3}\right)=0.
\end{equation}
The paraxial equation for a freely propagating field is obtained from equation (\ref{ErmaF0}). Its solution, equation (\ref{ermasol}), produces the values
\begin{equation}
 \nu=\frac{1}{2}\ln (az^2+bz +c),\qquad   \mu=-\frac{1}{2}\int \frac{dz}{az^2+bz+c},
\end{equation}
and the conditions $\nu(0)=0 \rightarrow c=1$ and  $a-b^2/2=1$. \textcolor{black}{The choice} $S(x,0)=0=Q(x)\dot{\nu}(0)+\mu(0)$ implies $\nu(0)=\mu(0)=0$, that delivers the values for $a$ and $b$ given by
\begin{equation}
    \dot{\nu}=\frac{1}{2}\frac{2az+b}{az^2+bz +c}, \qquad  \dot{\nu}(0)=0\rightarrow b=0 \, , a=1.
\end{equation}
Therefore, we may write finally the amplitude
\begin{equation}
A(x,z)=\frac{1}{\pi^{1/4}(z^2 +1)^{1/4}}\exp\left[-\frac{x^2}{2(z^2 +1)}\right],
\end{equation}
which is nothing but the amplitude for a Gaussian beam freely (paraxially) propagating. The phase is written as
\begin{equation}
    S(x,z)=\frac{z}{z^2+1}x^2-\frac{1}{2}\int \frac{dz}{z^2+1},
\end{equation}
that gives the Gouy phase
\begin{equation}
    \phi_G(z)=-\frac{1}{2}\arctan z.
\label{gouyphaseknown}
\end{equation}
In this way, we have obtained Gouy's phase from the Madelung-Bohm index of refraction, equation (\ref{0130}), where the quadratic term generates an effective GRIN medium \cite{ChileI} that in turn produces the focusing effect needed to induce the Gouy phase shift. 

\section{Conclusions}
We have used Ermakov-Lewis techniques to show that if we choose a (Bohm) phase that produces a Gaussian field, the Gouy phase arises naturally in the Madelung-Bohm approach to paraxial wave propagation. This is done with the help of the solution of the Ermakov equation coming from the explicitly time dependent invariant introduced by Lewis \cite{Lewis0,Lewis}, in our case for the Hamiltonian of the free particle (paraxial free propagation in classical optics).   \textcolor{black}{Although the Gouy phase \eqref{gouyphaseknown} has been  known for long time \cite{Gouy1,Gouy2} its physical origin is still a matter of research \cite{Michigan,Carmelo}. The procedure detailed here gives a natural explanation of its origin in terms of the form of the spread of the wave, which is measured by the Bohm index of refraction. It is expected that this gives hints on Gouy phases for different kinds of beams.}

\section{Conflict of interest}
The authors declare that they have no conflict of interest.

\section{Author contributions}
The authors declare to have contributed in the same way in the elaboration of this article.

\end{document}